# Patricia’s Bad Distributions

### v1

@@ 1. Introduction and results
PATRICIA trees have been analysed assuming various probabilistic models for the input strings; where usually the infinite strings are assumed to be independent and identically distributed over $\{0,1\}^{\mathbb{N}}$. Note that atoms of such a distribution result in identical strings with positive probability, and in this case the construction of the trie does not lead to a finite tree. Hence, the law of the strings is usually assumed to be diffuse (non-atomic). Special cases of such diffuse probability distributions have been considered in the analysis of algorithms on strings such as the Bernoulli models, Markov model, dynamical sources or the density model; see [5, 3, 13, 10, 18, 14, 12] and the references given in these papers.
Figure 1. On the left the trie for the strings $00000\ldots$, $00001\ldots$, $0100\ldots$, $0101\ldots$, $1100\ldots$, and $1101\ldots$ is shown. It’s leaves are the full black vertices, the indicated children of the full black vertices do not belong to the trie. Vertices with out-degree $1$ within the trie are indicated by arrows. On the right the resulting PATRICIA tree by deleting corresponding edges is shown.
In the present note we focus on the height of a PATRICIA tree, which is the maximal (graph) distance of any leave from the root. The asymptotic behavior of the height of tries under the Bernoulli models is covered in Devroye [4, 6]. For general diffuse laws concentration of the height of PATRICIA trees is studied (assuming only independence of the infinite strings not necessarily identical distribution) in Devroye [7] based on results from [1]; see also [15] for concentration of the height of PATRICIA trees in the Bernoulli model.
While such studies aim to show that the height behaves well with respect to applications from algorithms, Evans and Wakolbinger [8, 9] studied these random tree structures as tree-valued transient Markov chains from the perspective of Doob–Martin boundary theory. They asked (private communication) how high PATRICIA trees can grow for arbitrary diffuse probability distributions of the strings (see [9, Section 5] for specific examples). The subject of the present note is to answer this question by Theorems 1 and 2: The expected height grows always sublinearly, but can be made to exceed any fixed sublinear growth rate by the choice of an appropriate diffuse law.
@@
For any sequence $\alpha=(\alpha_{n})_{n\in{\mathbb{N}}}$ of positive numbers with $\alpha_{n}\to\infty$ as $n\to\infty$ there exists a diffuse probability distribution $\mu=\mu^{(\alpha)}$ on $\{0,1\}^{\mathbb{N}}$ such that for all $n$ sufficiently large
$\,{\mathbb{E}}[H^{\mu}_{n}]\geq\frac{n}{\alpha_{n}}.$
We call a law $\mu$ on $\{0,1\}^{\mathbb{N}}$ causing large expected heights $\,{\mathbb{E}}[H^{\mu}_{n}]$ bad since such laws are undesirable from the point of view of the efficiency of algorithms based on PATRICIA trees. The remaining part of the present note contains proofs of these two theorems.
@@ 2.1. Proof of Theorem 1
Suppose for a contradiction that there exists $\varepsilon>0$ such that for all $k\in{\mathbb{N}}$ there is a string $v_{1}\ldots v_{k}\in\{0,1\}^{k}$ such that ${\mathbb{P}}(\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k})\geq\varepsilon$. Then by a compactness argument shown below there exists an infinite string $v=(v_{i})_{i\in{\mathbb{N}}}\in\{0,1\}^{\infty}$ such that for all $k\in{\mathbb{N}}$, ${\mathbb{P}}(\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k})\geq\varepsilon$. The events $\{\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k}\}$ are decreasing in $k$, so this implies that
@@
It remains to show the existence of the infinite string $v=(v_{i})_{i\in{\mathbb{N}}}\in\{0,1\}^{\mathbb{N}}$ such that for all $k\in{\mathbb{N}}$, ${\mathbb{P}}(\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k})\geq\varepsilon$. Consider $\{0,1\}$ as a topological space with the discrete topology (all subsets being open) and $\{0,1\}^{\mathbb{N}}$ as the product space with the product topology. As a product of compact spaces $\{0,1\}^{\mathbb{N}}$ is compact, and the projections $\Pi_{k}:\{0,1\}^{\mathbb{N}}\to\{0,1\}^{k}$ given by
@@
is closed and thus compact in $\{0,1\}^{\mathbb{N}}$. This implies that $(V_{k})_{k\in{\mathbb{N}}}$ is a nested sequence of non-empty, compact sets. Hence, we have
@@ 2.2. Proof of Theorem 2
where $|A|$ denotes the cardinality of $A$, i.e., the number of distinct elements within $\{\langle\Theta^{(1)}\rangle,\ldots,\langle\Theta^{(n)}\rangle\}$. For all $1\leq i<j\leq n$ we have ${\mathbb{P}}(\langle\Theta^{(i)}\rangle=\langle\Theta^{(j)}\rangle)=1/N^{2}$. Hence, we obtain
@@
The probability distribution $\mu^{(\alpha)}$ on $\{0,1\}^{\mathbb{N}}$ is obtained in two stages. First, let $G$ be a random variable with geometric distribution with parameter $\frac{1}{2}$, i.e., with ${\mathbb{P}}(G=k)=(\frac{1}{2})^{k}$ for $k\in{\mathbb{N}}$. Then define a sequence $(\lambda_{i})_{i\in{\mathbb{N}}}$ by
@@
where $\Theta=(\vartheta_{i})_{i\in{\mathbb{N}}}$ is independent of $G$ and has law $\mu_{A(k)}$ defined in (2) with $A(\cdot)$ defined in (5). We then define $\mu=\mu^{(\alpha)}$ as the law of $\Lambda=(\lambda_{i})_{i\in{\mathbb{N}}}$. Since the $\mu_{A(k)}$ are diffuse, we obtain that $\mu$ is diffuse.
@@
is $\mathrm{Bin}(n,2^{-\beta_{n}})$-distributed. To get rid of the floors in the definition of $\beta_{n}$ denote by $X^{\prime}_{n}$ a $\mathrm{Bin}(n,4/\alpha_{n})$-distributed random variable. Note that $2^{-\beta_{n}}\geq 4/\alpha_{n}$. By Okamoto’s inequality, see [21] or [2, Exercise 2.12] we have
@@
$\geq\frac{n}{\alpha_{n}}.\qed$
@@ Acknowledgement
This research was mainly done during the Sixteenth Annual Workshop on Probability and Combinatorics at McGill University’s Bellairs Institute in Holetown, Barbados. We thank Bellairs Institute for its hospitality and support. We also thank Jasmin Straub for comments on a draft of this note.

### v2

@@ 1. Introduction and results
PATRICIA trees have been analysed assuming various probabilistic models for the input strings; where usually the infinite strings are assumed to be independent and identically distributed over $\{0,1\}^{\mathbb{N}}$. Note that atoms of such a distribution result in identical strings with positive probability, and in this case the construction of the trie does not lead to a finite tree. Hence, the law of the strings is usually assumed to be diffuse (non-atomic). Special cases of such diffuse probability distributions have been considered in the analysis of algorithms on strings such as the Bernoulli models, Markov model, dynamical sources or the density model; see [23, 24, 6, 4, 14, 11, 19, 1, 15, 13] and the references given in these papers.
Figure 1. On the left the trie for the strings $00000\ldots$, $00001\ldots$, $0100\ldots$, $0101\ldots$, $1100\ldots$, and $1101\ldots$ is shown. Its leaves are the full black vertices, the indicated children of the full black vertices do not belong to the trie. Vertices with out-degree $1$ within the trie are indicated by arrows. On the right the resulting PATRICIA tree by deleting corresponding edges is shown.
In the present note we focus on the height of a PATRICIA tree, which is the maximal (graph) distance of any leave from the root. The asymptotic behavior of the height of tries and PATRICIA trees under the Bernoulli models is covered by Pittel [23, 24] and Devroye [5, 7]. For example, for the height $H^{\mathrm{syB}}_{n}$ of the PATRICIA tree constructed from $n$ independent strings under the symmetric Bernoulli model, i.e. all bits being independent and Bernoulli$(\frac{1}{2})$ distributed, Pittel [23] obtained as $n\to\infty$ that
$\frac{H^{\mathrm{syB}}_{n}}{\log n}\to 1\mbox{ almost surely.}$.
This shows an asymptotic 50% improvement of the PATRICIA tree over the trie, for which the limit constant for the same probabilistic model is $2$ instead of $1$. For general diffuse laws concentration of the height of PATRICIA trees is studied (assuming only independence of the infinite strings not necessarily identical distribution) by Devroye [8] based on results from [2]; see also [16] for concentration of the height of PATRICIA trees in the Bernoulli model.
While such studies aim to show that the height behaves well with respect to applications from algorithms, Evans and Wakolbinger [9, 10] studied these random tree structures as tree-valued transient Markov chains from the perspective of Doob–Martin boundary theory. They asked (private communication) how high PATRICIA trees can grow for arbitrary diffuse probability distributions of the strings (see [10, Section 5] for specific examples). The subject of the present note is to answer this question by Theorems 1 and 2: The expected height grows always sublinearly, but can be made to exceed any fixed sublinear growth rate by the choice of an appropriate diffuse law.
@@
For any sequence $\alpha=(\alpha_{n})_{n\in{\mathbb{N}}}$ of positive numbers with $\alpha_{n}\to\infty$ as $n\to\infty$ there exists a diffuse probability distribution $\nu=\nu^{(\alpha)}$ on $\{0,1\}^{\mathbb{N}}$ such that for all $n$ sufficiently large
$\frac{\,{\mathbb{E}}[H^{\nu}_{n}]}{n/\alpha_{n}}\to\infty,\quad\mbox{ and }\quad\frac{H^{\nu}_{n}}{n/\alpha_{n}}\to\infty\mbox{ almost surely.}$.
We call a law $\nu$ on $\{0,1\}^{\mathbb{N}}$ causing large expected heights $\,{\mathbb{E}}[H^{\nu}_{n}]$ bad since such laws are undesirable from the point of view of the efficiency of algorithms based on PATRICIA trees. The remaining part of the present note contains proofs of these two theorems.
Remark 3.
For the density model, which is a subclass of the diffuse distributions on $\{0,1\}^{\mathbb{N}}$, the asymptotics of Theorem 1 were obtained by Devroye [6, page 419]. There, also bad distributions with asymptotic properties as in our Theorem 2 are constructed for sequences $\alpha_{n}=n^{\varepsilon}$ with $10<\varepsilon<$ 1.
@@ 2.1. Proof of Theorem 1
Suppose for a contradiction that there exists $\varepsilon>0$ such that for all $k\in{\mathbb{N}}$ there is a string $v_{1}\ldots v_{k}\in\{0,1\}^{k}$ such that ${\mathbb{P}}(\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k})\geq\varepsilon$. Then by a compactness argument shown below there exists an infinite string $v=(v_{i})_{i\in{\mathbb{N}}}\in\{0,1\}^{{\mathbb{N}}}$ such that for all $k\in{\mathbb{N}}$, ${\mathbb{P}}(\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k})\geq\varepsilon$. The events $\{\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k}\}$ are decreasing in $k$, so this implies that
@@
It remains to show the existence of the infinite string $v=(v_{i})_{i\in{\mathbb{N}}}\in\{0,1\}^{\mathbb{N}}$ such that for all $k\in{\mathbb{N}}$, ${\mathbb{P}}(\xi_{1}\ldots\xi_{k}=v_{1}\ldots v_{k})\geq\varepsilon$. Consider $\{0,1\}$ as a topological space with the discrete topology (all subsets being open) and $\{0,1\}^{\mathbb{N}}$ as the product space with the product topology. As a product of compact spaces $\{0,1\}^{\mathbb{N}}$ is compact. It is also a Hausdorff space. The projections $\Pi_{k}:\{0,1\}^{\mathbb{N}}\to\{0,1\}^{k}$ given by
@@
is closed and thus compact in $\{0,1\}^{\mathbb{N}}$. This implies that $(V_{k})_{k\in{\mathbb{N}}}$ is a nested sequence of non-empty, compact sets. Now, Cantor’s intersection theorem implies
@@ 2.2. Proof of Theorem 2
where $|A|$ denotes the cardinality of $A$, i.e., the number of distinct elements within the set $\{\langle\Theta^{(1)}\rangle,\ldots,\langle\Theta^{(n)}\rangle\}$. For all $1\leq i<j\leq n$ we have ${\mathbb{P}}(\langle\Theta^{(i)}\rangle=\langle\Theta^{(j)}\rangle)=1/N^{2}$. Hence, we obtain
@@
First, a probability distribution $\mu^{(\alpha)}$ on $\{0,1\}^{\mathbb{N}}$ is obtained in two stages. Let $G$ be a random variable with geometric distribution with parameter $\frac{1}{2}$, i.e., with ${\mathbb{P}}(G=k)=(\frac{1}{2})^{k}$ for $k\in{\mathbb{N}}$. Then define a sequence $(\lambda_{i})_{i\in{\mathbb{N}}}$ by
@@
where $\Theta=(\vartheta_{i})_{i\in{\mathbb{N}}}$, conditional on $\{G=k\}$, has law $\mu_{A(k)}$ defined in (2) with $A(\cdot)$ defined in (5). We then define $\mu=\mu^{(\alpha)}$ as the law of $\Lambda=(\lambda_{i})_{i\in{\mathbb{N}}}$. Since the $\mu_{A(k)}$ are diffuse, we obtain that $\mu$ is diffuse.
@@
is $\mathrm{Bin}(n,2^{-\beta_{n}})$-distributed. To get rid of the floors in the definition of $\beta_{n}$ denote by $X^{\prime}_{n}$ a $\mathrm{Bin}(n,4/\alpha_{n})$-distributed random variable. Note that $2^{-\beta_{n}}\geq 4/\alpha_{n}$. By Okamoto’s inequality, see [22] or [3, Exercise 2.12] we have
@@
$\geq\frac{n}{\alpha_{n}}.$
Since the sequence $(\log\alpha_{n})$ tends to infinity the present proof implies the existence of a diffuse probability distribution $\nu=\nu^{(\alpha)}$ on $\{0,1\}^{\mathbb{N}}$ such that $\,{\mathbb{E}}[H^{\nu}_{n}]\geq n/\log\alpha_{n}$ for all sufficiently large $n\in{\mathbb{N}}$, hence
$\frac{\,{\mathbb{E}}[H^{\nu}_{n}]}{n/\alpha_{n}}\to\infty.$
To prove the second statement of Theorem 2 we use the following bound from Devroye [8, page 21]: for any diffuse probability distribution $\mu$ on $\{0,1\}^{\mathbb{N}}$ and any $t>0$,
(10) ${\mathbb{P}}\left(H^{\mu}_{n}\leq\,{\mathbb{E}}[H^{\mu}_{n}]-t\right)\leq\exp\left(-\frac{t^{2}}{2\,{\mathbb{E}}[H^{\mu}_{n}+1]}\right)\leq\exp\left(-\frac{t^{2}}{2n}\right).$
We now consider the probabilities
(11) ${\mathbb{P}}\left(H^{\nu}_{n}\leq\frac{n}{\log^{2}\alpha_{n}}\right)={\mathbb{P}}\left(H^{\nu}_{n}\leq\,{\mathbb{E}}[H^{\nu}_{n}]-\left(\,{\mathbb{E}}[H^{\nu}_{n}]-\frac{n}{\log^{2}\alpha_{n}}\right)\right)$
and note that for all sufficiently large $n$ we have
(12) $\,{\mathbb{E}}[H^{\nu}_{n}]-\frac{n}{\log^{2}\alpha_{n}}\geq\frac{n}{\log\alpha_{n}}-\frac{n}{\log^{2}\alpha_{n}}=n\,\frac{\log(\alpha_{n})-1}{\log^{2}\alpha_{n}}\geq\frac{n}{\log^{2}n}.$
Combining (10)–(12) we obtain
${\mathbb{P}}\left(H^{\nu}_{n}\leq\frac{n}{\log^{2}\alpha_{n}}\right)$ $\leq\exp\left(-\frac{n}{2\log^{4}n}\right)$
for all sufficiently large $n$. Since these upper bounds are summable it follows from the first Borel–Cantelli Lemma that $\liminf_{n\to\infty}H^{\nu}_{n}/(n/\log^{2}\alpha_{n})\geq 1$ almost surely, hence
$\frac{H_{n}^{\nu}}{n/\alpha_{n}}\to\infty\mbox{ almost surely.}$.
Thus, $\nu$ has the properties claimed in Theorem 2. ∎
@@ Acknowledgement
This research was mainly done during the Sixteenth Annual Workshop on Probability and Combinatorics at McGill University’s Bellairs Institute in Holetown, Barbados. We thank Bellairs Institute for its hospitality and support. We also thank the referees and Jasmin Straub for comments on a draft of this note.